From nch@roe.ac.uk Fri May 6 11:03:27 2005 Date: Fri, 6 May 2005 10:57:31 +0100 (BST) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Mike Read , Nigel Hambly , Nicholas Cross , Bob Mann Cc: CCs for WSA weekly meeting minutes distribution -- Andrew Lawrence , Andy Adamson , John Taylor , Jim Emerson , Malcolm Stewart , Martin Hill , Mike Irwin , Peredur Williams , Stephen Warren Subject: WFAU WSA weekly meeting minutes, 6th May 2005 Minutes of WFCAM Science Archive meeting: 6th May 2005 ------------------------------------------------------------- ------------------------------------------------------------- Present: NCH, NJC, PMW, RGM, MAR, ETWS Apologies: MCH, AL, JMS, JPE DONM: 10am, Friday 13th May 2005, plate library Actions discharged: ------------------- ACTION: ETWS to look into 3dm RAID monitoring on sneferu, along the lines of the browsable pages on djoser. Discharged; 3dm logs now browsable on sneferu as with djoser. NCH & ETWS have now agreed that ETWS will routinely monitor the 3dm SATA RAID arrays while NCH will monitor the LSI SCSI arrays on the W2K3 systems (until the archive operator post is filled). Actions partly discharged but continuing: ----------------------------------------- None this week. Actions carried forward from 15/04/05 meeting: ---------------------------------------------- ACTION: NCH to write a tutorial on how to query cross-matched data in the WSA for the web pages. - CONTINUES Specific points and new actions: -------------------------------- Project management: VDMT meeting (attended by PMW, NCH and JMS) passed off without major issues yesterday. PMW noted that the paperwork for the archive operator position has gone to UoE HR; RGM noted that he and PMW need to put the finishing touches to the datamining post particulars before sending them off. WFCAM update: Nothing new to report; check linked pages on the TWiki for up-to-date information on status of SV and commencement of surveys proper. Comments and issues arising from CASU fortnightly minutes: Nothing new this week. Networking: Nothing new this week. Hardware: NCH noted that everything continues to work normally with the exception of a power outage last night; good news is that the online public DB server amenhotep survived, while the offline load server ahmose has failed to reboot (helpdesk submitted to IT support to look into this; looks like a RAID array problem). Software: ETWS & NCH have conintued to work hard on CUs 4 & 7 (enhancements to CU4 include full precision RA/Dec computations for ingest; CU7 is being enhanced for merged source classification - thanks to MJI for a cunning Bayesian scheme for this - along with a few new attributes to help users weed out spurious pairings, and testing of seaming across large area surveys). MAR has completed implementation of web applications for community administration - these are now being looked at by the UKIDSS bosses SJW and AL. Finally, NJC noted that his first draft of the global photometric calibration document is nearly done. ACTION: NJC to distribute the first draft of the global photometric calibration document by close of play today (Friday 6th May). Data Release: No further news this week SSA: Nothing new this week Astrogrid deployment: No news this week. Miscellaneous: RGM reminded the team about the meeting concerning experimentation with the University's Storage Area Network and high-end database servers: Mon 9th at 9:30, New Building meeting room #2 (upstairs); attendees RGM, PMW and NCH.